An investor holds 5 000 shares of VC Ltd, which was bought for R18 per share and is now priced at R20 per share. VC Ltd makes a 2-for-3 rights issue at R15.80. The investor decides to subscribe to the rights issue and purchases a further 3 000 shares. What is the price per share expected to adjust to post the rights issue (the ex-rights price)?

R 18.43

R 17.38

R 17.90

R17.18
Answer #1

First let us find out the number of right shares investor can subscribe:

2-for-3 rights issue.

That means number of right-shares investor can subscribe = (5000 / 3) * 2 = 3,333 shares.

Therefore, investor will get that 3000 shares at R$15.80

The amount to be paid = number of shares subscribed * Price of right-share

= 3000 * 15.80 = R$47,400

The amount paid to get previous 5000 shares = 5000 * R$18 = R$90,000

Therefore, total amount incurred to buy 8000 shares = R$90,000 + R$47,400 = R$137,400

Therefore, adjusted price per share post issue = Total amount incurred to buy 8000 shares / total number of shares.

= R$137,400 / 8000 = R$17.18